OpenStudy (anonymous):

Jasmin traveled a certain distance at exactly 35 mph and then she traveled another distance at exactly 55 mph. Write an equation in standard form that models this situation if Jasmin went a total distance of 620 miles.

jimthompson5910 (jim_thompson5910):

amishjeb, the formula is d = r*t not t = r*d

OpenStudy (anonymous):

True.

jimthompson5910 (jim_thompson5910):

if you let x = time spent going 35 mph y = time spent going 55 mph note: both times are in hours then 35x is the distance you go 35 mph for x hours 55y is the distance you go 55 mph for y hours

OpenStudy (anonymous):

@jim_thompson5910 so it would look like 620=35+55t?

jimthompson5910 (jim_thompson5910):

add the two distances to get 35x+55y the total distance traveled is 35x+55y and this is for x+y hours

jimthompson5910 (jim_thompson5910):

we're told the total distance is 620 miles, so this means 35x+55y = 620

OpenStudy (anonymous):

Jim, I had something else come up like this. Maybe you could help me. Can you only use one variable and set it up like this: 35x +55(620-x)=620?

jimthompson5910 (jim_thompson5910):

you could rewrite it, but that's effectively what they want

jimthompson5910 (jim_thompson5910):

no because x+y is the total TIME, not distance

jimthompson5910 (jim_thompson5910):

you cannot say x+y = 620 and then rewrite it to get y = 620 - x

jimthompson5910 (jim_thompson5910):

if you knew that it took 10 hours, then you could say x+y = 10 ---> y = 10 - x but we don't have that kind of info

OpenStudy (anonymous):

Ahhhh alright. Thanks. I'll try not to post bad info next time :) Sorry guys.

jimthompson5910 (jim_thompson5910):

that's ok, it's all part of the learning process

jimthompson5910 (jim_thompson5910):

don't be afraid to get it wrong

OpenStudy (anonymous):

so basically you sub. in ax+by=c for a=35 and b=55 and c= 620 all together

jimthompson5910 (jim_thompson5910):

yes using the steps I showed above and you could reduce things, but 35x+55y = 620 is an equation in standard form, so you can stop there

jimthompson5910 (jim_thompson5910):

you could keep going and divide everything by 5 to get 35x/5 = 7x 55y/5 = 11y 620/5 = 124 so 35x+55y = 620 turns into 7x + 11y = 124 but this step is optional really

OpenStudy (anonymous):

y would you divide everything by 5 im confused

jimthompson5910 (jim_thompson5910):

5 is the GCF of the numbers 35, 55, 620

jimthompson5910 (jim_thompson5910):

you could divide everything by the GCF so it fully reduces (and everything will stay a whole number)

OpenStudy (anonymous):

oh ohk so you could do the same for You are going shopping for some new school clothes. You see some shirts and jeans that you like. The shirts are $35.99 each and the jeans are $65.00 each. You have $350 to spend.

OpenStudy (anonymous):

so it would be 35.99x+65y=350 and then the gcf =5 ? right

jimthompson5910 (jim_thompson5910):

in this case, the GCF wouldn't apply because 35.99 isn't a whole number

jimthompson5910 (jim_thompson5910):

what you could do though is multiply everything by 100 to make 35.99 turn into 3599 (a whole number) 35.99x*100 = 3599x 65y*100 = 6500y 350*100 = 35000 so the equation 35.99x+65y=350 becomes 3599x+6500y = 35000 which is now in standard form

OpenStudy (anonymous):

why would you multiply by 100

jimthompson5910 (jim_thompson5910):

because standard form is Ax+By = C where A,B,C are whole numbers (A is usually nonnegative)

OpenStudy (anonymous):

but why 100 though

jimthompson5910 (jim_thompson5910):

well to move the decimal point 2 spots to the right, you multiply by 100

jimthompson5910 (jim_thompson5910):

if we had 3.599 and we wanted to move the decimal 3 spots to the right, we'd multiply it by 1000

OpenStudy (anonymous):

oh ok thank you for breaking it down it really makes alot of sense now

jimthompson5910 (jim_thompson5910):

I'm glad it's all clicking, you're welcome

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@jim_thompson5910 i have one more question